# ALERT-Transformer: Bridging Asynchronous and Synchronous Machine Learning for Real-Time Event-based Spatio-Temporal Data

### v2

@@ 1 Introduction
Event-based sensors capture visual information in an event-driven, asynchronous manner (Finateu et al., 2020; Gallego et al., 2020). Efficiently exploiting their data has proven challenging as the vast majority of approaches published in the literature consist of either converting event-based data to dense representations, or deploying spiking neural networks (SNNs) on streams of events. The former allows to exploit standard machine learning (ML) frameworks such as PyTorch and Tensorflow, but does not leverage the inherent sparsity and other properties of event-based data (Gehrig et al., 2019). The latter relies on SNNs, which are hard to train and usually exhibit lower accuracy than an equivalent dense neural network. Furthermore, while the neuromorphic community has argued in favor of their higher energy efficiency for decades, recent research and breakthroughs in edge AI accelerators suggest that this might not be the case (Dampfhoffer et al., 2023; Garrett et al., 2023; Moosmann et al., 2023; Caccavella et al., 2023).
@@
Standard ML relies on tensor-based processing. Converting the stream of events –represented as tuples of values (x and y coordinates, polarities and timestamp) – to a multidimensional tensor is thus a crucial step. The challenge involves (1) representing time in a reliable and continuous manner, allowing it to be processed similarly to the finite spatial and polarity dimensions, (2) continuously incorporating new events in the feature tensors which also requires forgetting previous events, (3) using limited computational resources to allow real-time processing. Our main contributions towards Event-Based ML are the following:
  • We introduce a system that allows converting an asynchronous stream of events into a synchronous list of feature vectors.
  • We propose a solution to represent continuous time as bi-dimensional vector of bounded values, inducing negligible decrease of 00 to $0.4$% relative accuracy.
  • We propose a method to hybridize a PointNet model into an asynchronous version, without modifying its internal neural modules, by adding a decay mechanism applied on the output embeddings.
  • We demonstrate that the resulting embedding can be sampled at any arbitrary time, down to a per-event operation, allowing ultra-low latency decision making.
  • Finally, we present a framework and associated ALERT-Transformer model, to train a network on event-based data end-to-end, and then operate it in synchronous regime for high accuracy on gesture recognition, or asynchronously for ultra-low latency.
Our idea is that, even if the sensor outputs data at 1MHz or more (Finateu et al., 2020), the application/client side will not, and does not need to, run at this speed. For instance, in a embedded product, such as a drone, that relies on event-based sensors for self-localization in space, this localization step would not operate at more than 100Hz (Kaufmann et al., 2023). Therefore, we consider that this hybrid (asynchronous-to-synchronous) conversion module would be operated as close to the sensor as possible, inside a 3D-stacked image sensor, as proposed in (Bouvier et al., 2021; Bonazzi et al., 2023).
@@ 1.1.1 Event-Based Data Representations
Event-based data is extremely sparse, with unpredictable sparsity patterns, making it unsuitable for processing raw with dense machine learning pipelines. An extended approach consists in integrating events over fixed time windows, creating frame-like representations using methods such as histograms or event queues (Innocenti et al., 2021; Sabater et al., 2022; Maqueda et al., 2018; Zhu et al., 2018). To maximize task accuracy from end-to-end, (Gehrig et al., 2019) proposed learning the kernel for convolving the stream of events into a discreet tensor, called “Event-Spike Tensor”. While the resulting networks perform better, this method still requires waiting to accumulate all events before starting processing and is thus not asynchronous.
@@ 1.1.2 PointNet Architecture
The PointNet (Qi et al., 2017a, b) is an architecture designed for processing three-dimensional (3D) point clouds. It extracts features from individual points with a shared Multi-Layer Perceptron (MLP). It then reduces the entire point cloud with a max-pooling operation into a single global feature vector. Given the similarity in data representation, several works interpreted event streams as point clouds (Wang et al., 2019a; Chen et al., 2022; Zhao et al., 2021; Wang et al., 2019b). These approaches mostly treat time as any other coordinate, making fast and efficient processing challenging.
EventNet (Sekikawa et al., 2019) addresses this issue by processing time separately from other coordinates. The authors propose a modification to the $max$ operator which becomes a recursive function capable of updating temporal and spatial information as new events arrive. Our solution, introduced in Section 2, builds upon theirs, but eliminates the need to modify the $max$ operator. The structure of the PointNet remains unaltered and is thus entirely compatible with existing AI accelerators.
@@ 1.1.3 Event-Based Transformers
Our work takes inspiration from PointBERT (Yu et al., 2022), which combines PointNets with a Transformer (Vaswani et al., 2017) for 3D classification. Models proposed by (Sabater et al., 2022, 2023; Wang et al., 2022; Peng et al., 2023) generate frames of aggregated events using various strategies, and convert them into tokens with a patch-to-token strategy inspired by the Vision Transformer (ViT [^1] In ViT, images are split into fixed-size patches, which are then linearly embedded to obtain a sequence of vectors.) (Dosovitskiy et al., 2021). They leverage spatial sparsity by discarding input patches lacking sufficient events. In (Peng et al., 2023), grouped convolutions are used to embed successive patches through time, each token thus handles information from different time steps. (Chen et al., 2022) attempts to input a reduced 3D event cloud to the Point Transformer (Zhao et al., 2021), a model originally designed for spatial 3D point clouds. All these approaches demonstrate high accuracy, motivating our choice of using Transformer models. However, most are not trained end-to-end, and none can process events asynchronously.
@@ 1.1.4 Asynchronous Processing
Beside (Sekikawa et al., 2019), some works focus on asynchronous processing of event streams with standard network architectures. (Sironi et al., 2018) presented an architecture using local memory units shared by neighboring pixels. (Messikommer et al., 2020) introduced Asynchronous Sparse Convolutional Networks, a framework for converting models trained on synchronous image-like event representations into asynchronous models. Even though these solutions are built for event-driven processing, they add on complexity with respect to standard ML tools.
@@ 2 Proposed Architecture
Figure 2: Overview of a full (A)LERT-Transformer model. The event stream is spatially divided into local event clouds. The red crosses indicate the non-active tokens based on their number of events. An (A)LERT module converts them to individual high dimension features (tokens), which are fed to a Transformer and classifier head.
@@ 2.1.1 Raw Events to Normalized Local Patches
The embedding part – which we call Learnt Embedding for Real-Time Processing of Event-Based Data (LERT) Module – is built upon the PointNet architecture (Qi et al., 2017a). LERT spatially divides the input stream in local event clouds (patches) and converts each of them into individual feature vectors (a.k.a. tokens). This module is depicted in Figure 3. To extract patches from an event sequence, a grid in the (x, y)-plane is constructed, effectively dividing the space into same-sized groups of pixels with predefined coordinates. A local event cloud is simply a set of events, which is a subset of the original input event stream.
Once the events are organized in patches, the LERT modules applies two transformations. First, all patches which do not contain enough events are filtered out. The filter mechanism is a simple threshold on the number of events present inside the local point cloud. The threshold value is fixed and searched via a hyperparameter search. It could be trained. We denote the removed patches and events they contain as non-active. This method was proposed by (Sabater et al., 2022) and allows to exploit the sparsity of the input event-based data. The remaining events and patches are referred to as active events and active patches. The second step is a trivial normalization, where the spatial coordinates $(x,y)$ of the active events are scaled down to the range $[-1,1]$ with respect to a patch size (not the full image size). Once this preprocessing of the events is finished, the active event coordinates inside each patch are represented as follows:
@@
$T$ denotes the duration of the input event stream sample (not full file duration). LERT operates in a fully synchronous manner which makes it compatible with widespread machine learning tools and frameworks, which simplifies training. It takes as input a finite list of events whose length ($Ne$) depends on the event accumulation mode. $Ne$ is either always the same (we call it Constant Count Input Mode (CCIM)) or varying and depending on the number of events triggered during a fixed time window (denoted Constant Time Input Mode (CTIM)), in which case $T$ is always the same between two samples.
@@ 2.1.2 Normalized Local Patches to Tokens
The Feature Generator (FG) then converts each normalized input patch into an input token for the transformer model. The FG is shared between all the patches and consists in a customized shallow PointNet architecture. Every 4D active event contained in a single patch is processed through a shared MLP unit to generate a single high dimensional vector. The feature generator inherits the capacity of the PointNet to process an input of varying length. In our implementation, an MLP is composed of a succession of 1D-convolution layers.
@@
Each 4D event is thus mapped to a high dimensional vector, and the Patch Feature is obtained by applying a channel-wise maximum operation on the resulting vectors. At this stage, all $P$ active patches have been converted into $P$ $c$-dimensional vectors, denoted as Patch Features (see Figure 4 (a)).
The LERT module then adds a positional embedding to each Patch Feature, which provides information about its spatial location within the predefined 2D grid. This positional embedding is learnt during training, and is unique for every patch coordinate. A detailed schematic of the LERT embedding can be seen Figure 3.
@@ 2.2.1 Introduction to the ALERT module
We thus introduce the Asynchronous LERT (ALERT) Module. This variation is designed to efficiently update the token sequence as new events occur (asynchronous processing), making it particularly suited for online inference of event-based data. To achieve this, the ALERT module requires three key modifications: handling continuous values of time, updating tokens with each new event, and memorizing past information.
@@ 2.2.2 Handling Continuous Time Values
The LERT module alters the time coordinates of events in a sequence, shifting each input stream so that the first event starts at time 0. This approach introduces inefficiencies in real-time asynchronous processing, as it hinders the feasibility of recurrent or iterative processing of events. Our proposed solution consists in encoding the time value by using a periodic function. We achieve this by representing time into two sinusoidal waves: $t_{x}(t)=\alpha\cos(2\pi ft+\phi)$, and $t_{y}(t)=\alpha\sin(2\pi ft+\phi)$; taking inspiration from the positinal encoding originally used in Transformers. The amplitude ($\alpha$), frequency ($f$) and phase ($\phi$) of these waves are constant and tuned for the application. The Time Encoded (TE) LERT (TELERT) Module applies this periodical wrap prior to the PointNet model. It is the only change made before the channel-wise maximum operation between the LERT and ALERT modules (see Figure 4 b). Next, we present a method applied at token stage that emulates a sliding buffer in input.
@@ 2.2.3 Asynchronous Event-Based Token Update
Add Information: The new event(s) is processed by TELERT and the resulting token is the channel-wise maximum between the previous token and the event feature vector.
Forgetting information using Old Maximum Value Decay (OMVD): All tokens’ channels whose associated counter channels are higher than a certain threshold undergo decay. All tokens are subject to this decay, not only the newly updated one, as the oldest information might be stored in different tokens. The “age” of the last update of each channel is simply tracked by a counter. The decay is applied in the form of an exponential (×$\times\text{e}^{-\lambda}$), where the decay rate $\lambda$ can be tuned for each model or scene characteristics, and could potentially be learned; the pseudocode is in Appendix A.1. This mechanism draws inspiration from the “leakage” concept in SNNs literature (Bouvier et al., 2019), but instead of applying it to all neurons/features of the model, we deploy it on the tokens only.
@@ 2.2.4 Token Memorization and Activity Status
In ALERT, each new event triggers a corresponding patch token update. No raw event needs to be buffered. Instead, all tokens and associated age counter values are memorized throughout the entire runtime of the model, their patch being active or not. The non-activeness of the tokens is tracked using an event count for each patch. This counter is updated ($+1$ or $-1$) every time the token is updated (added information or OMVD).
@@ 2.3.1 Transformer Encoder for Classification
To evaluate the proposed LERT and ALERT modules, we plug them to a Vision Transformer model trained from scratch for classification tasks. As our focus is to assess the functionality of the proposed learnt synchronous and asynchronous embeddings modules, we deploy a stack of standard transformer encoders, as in the original ViT (Dosovitskiy et al., 2021). The final classification head simply consists in a single linear layer followed by a SoftMax activation function.
@@ 2.3.2 Synchronous Training and Asynchronous Inference
During training and for synchronous inference (using TELERT), all samples are processed independently from each other. So, classic GPU/TPU accelerated processing is done as usual. Once TELERT is trained, the embedding module can be converted to ALERT for inference by adding the event update functionalities (adding information and OMVD).
The ALERT version, used only for inference, can process events continuously. However, applying the feature generator on every single event without distributed processing is extremely time consuming. Hence, for simulation purposes, the asynchronous inference pipeline is assessed by updating and processing the tokens on-demand every $\Delta t$. This has no impact on the resulting accuracy, as the mathematics of the asynchronous update is ultimately not altered. Proof can be found in Appendix A.2.

### v3

@@ 1 Introduction
Event-based sensors capture visual information in an event-driven, asynchronous manner (Finateu et al., 2020; Gallego et al., 2020). Efficiently exploiting their data has proven challenging as the vast majority of approaches published in the literature consist of either converting event-based data to dense representations, or deploying spiking neural networks (SNNs) on streams of events. The former allows to exploit standard machine learning (ML) frameworks such as PyTorch and Tensorflow, but does not leverage the inherent sparsity and other properties of event-based data (Gehrig et al., 2019). The latter relies on SNNs, which are hard to train and usually exhibit lower accuracy than an equivalent dense neural network. Furthermore, while the neuromorphic community has argued in favor of their higher energy efficiency for decades, recent research and breakthroughs in edge AI accelerators indicate this is still an open question (Dampfhoffer et al., 2023; Garrett et al., 2023; Moosmann et al., 2023; Caccavella et al., 2023).
@@
Standard ML relies on tensor-based processing. Converting the stream of events – represented as tuples of values ($x$ and $y$ pixel coordinates, polarities and timestamp) – to a multidimensional tensor is thus a crucial step. The challenge involves (1) representing time in a reliable and continuous manner, allowing it to be processed similarly to the finite spatial and polarity dimensions, (2) continuously incorporating new events in the feature tensors which also requires forgetting previous events, (3) using limited computational resources to allow real-time processing. Our main contributions towards Event-Based ML are the following:
  • The ALERT module, an embedding based on PointNet which continuously integrates new events dismissing old ones via a leakage mechanism. This module introduces novel asynchronous embedding updates.
  • A flexible readout of the embedded data that can feed any downstream model with up-to-date features at different sampling rates, down to a per-event operation, allowing ultra-low latency decision making.
  • A patch-based approach inspired by Vision Transformer to exploit input sparsity and optimize efficiency.
  • A time encoding solution to represent continuous time as a bi-dimensional vector of bounded values, at the cost of negligible decrease of relative accuracy.
  • The ALERT-Transformer, a framework incorporating all of the above, which is trained on event-based data end-to-end. The model can then operate in synchronous regime for high accuracy on gesture recognition, or asynchronously for ultra-low latency.
Our idea is that, even if the sensor outputs data at 1MHz or more (Finateu et al., 2020), the application/client side will not, and does not need to, run at this speed. For instance, in an embedded product, such as a drone, that relies on event-based sensors for self-localization in space, this localization step would not operate at more than 100Hz (Kaufmann et al., 2023). Therefore, we consider that this hybrid (asynchronous-to-synchronous) conversion module would be operated as close to the sensor as possible, inside a 3D-stacked image sensor, as proposed in (Bouvier et al., 2021; Bonazzi et al., 2023).
@@ 1.1.1 Event-Based Data Representations
Event-based data is extremely sparse, with unpredictable sparsity patterns, making it unsuitable for raw processing with dense machine learning pipelines. An extended approach consists in integrating events over fixed time windows, creating frame-like representations using methods such as histograms or event queues (Innocenti et al., 2021; Sabater et al., 2022; Maqueda et al., 2018; Zhu et al., 2018). To maximize task accuracy from end-to-end, (Gehrig et al., 2019) proposed learning the kernel for convolving the stream of events into a discrete tensor, called “Event-Spike Tensor”. While the resulting networks perform better, this method still requires waiting to accumulate all events before starting processing and is thus not asynchronous.
@@ 1.1.2 PointNet Architecture
The PointNet (Qi et al., 2017a, b) is an architecture designed for processing three-dimensional point clouds. It extracts features from individual points with a shared Multi-Layer Perceptron (MLP). It then reduces the entire point cloud with a max-pooling operation into a single global feature vector. Given the similarity in data representation, several works interpreted event streams as point clouds (Wang et al., 2019a; Chen et al., 2022; Zhao et al., 2021; Wang et al., 2019b). These approaches mostly treat time as any other coordinate, making fast and efficient processing challenging.
EventNet (Sekikawa et al., 2019) addresses this issue by processing time separately from other coordinates. They propose a modification to the $max$ operator which becomes a recursive function capable of updating temporal and spatial information as new events arrive. Our solution, introduced in Section 2, eliminates the need to modify the $max$ operator. The structure of the PointNet remains unaltered and is thus entirely compatible with existing AI accelerators.
@@ 1.1.3 Event-Based Transformers
Our work takes inspiration from PointBERT (Yu et al., 2022), which combines PointNets with a Transformer (Vaswani et al., 2017) for 3D classification. Models proposed by (Sabater et al., 2022, 2023; Wang et al., 2022; Peng et al., 2023) generate frames of aggregated events using various strategies, and convert them into tokens with a patch-to-token strategy inspired by the Vision Transformer (ViT [^1] In ViT, images are split into fixed-size patches, which are then linearly embedded to obtain a sequence of vectors.) (Dosovitskiy et al., 2021). They leverage spatial sparsity by discarding input patches lacking sufficient events. In (Peng et al., 2023), grouped convolutions are used to embed successive patches through time, each token thus handles information from different time steps. (Chen et al., 2022) attempts to input a reduced 3D event cloud to the Point Transformer (Zhao et al., 2021), a model originally designed for spatial 3D point clouds. Following a different approach, (Blegiers et al., 2023) employ a Video Transformer Network using event-frames. All these approaches demonstrate high accuracy, motivating our choice of using Transformer models. However, most are not trained end-to-end, and none can process events asynchronously.
@@ 1.1.4 Asynchronous Processing
Besides (Sekikawa et al., 2019), some works focus on asynchronous processing of event streams with standard network architectures. (Sironi et al., 2018) presented an architecture using local memory units shared by neighboring pixels. (Messikommer et al., 2020) introduced Asynchronous Sparse Convolutional Networks, a framework for converting models trained on synchronous image-like event representations into asynchronous models. Even though these solutions are built for event-driven processing, they add on complexity with respect to standard ML tools.
@@ 2 Proposed Architecture
Figure 2: Overview of the (A)LERT-Transformer model. The event stream is spatially divided into local event clouds. The red crosses indicate the non-active tokens based on their number of events. (A)LERT module converts them to individual high dimension features, which are fed to a Transformer and classifier head.
@@ 2.1.1 Raw Events to Normalized Local Patches
The embedding part – which we call Learnt Embedding for Real-Time Processing of Event-Based Data (LERT) Module – is built upon the PointNet architecture (Qi et al., 2017a). LERT spatially divides the input stream in local event clouds (patches) and converts each of them into individual feature vectors (a.k.a. tokens). This module is depicted in Figure 4. To extract patches from an event sequence, a grid in the (x, y)-plane is constructed, effectively dividing the space into same-sized groups of pixels with predefined coordinates. Events triggered in the same pixel-group make a patch. Hence, a local event cloud is simply a set of events, which is a subset of the original input event stream.
Once the events are organized in patches, the LERT module applies two transformations. First, all patches which do not contain enough events are filtered out. The filter mechanism is a simple threshold on the number of events present inside the local point cloud. The threshold value is fixed and selected via a hyperparameter search. It could be trained. We denote the removed patches and events they contain as non-active. This method was proposed by (Sabater et al., 2022) and allows to exploit the sparsity of the input event-based data. The remaining events and patches are referred to as active events and active patches. The second step is a trivial normalization, where the spatial coordinates $(x,y)$ of the active events are scaled down to the range $[-1,1]$ with respect to a patch size (not the full image size). Once this preprocessing of the events is finished, the active event coordinates inside each patch are represented as follows:
@@
$T$ denotes the duration of the input event stream sample (not the full file duration). LERT operates in a fully synchronous manner which makes it compatible with widespread machine learning tools and frameworks, simplifying training. It takes as input a finite list of events whose length ($Ne$) depends on the event accumulation mode. $Ne$ is either always the same (we call it Constant Count Input Mode (CCIM)), or varying and depending on the number of events triggered during a fixed time window (denoted Constant Time Input Mode (CTIM)), in which case $T$ is always the same between two samples. Figure 3 provides a visualization of event data and the two possible input modes.
Figure 3: Schematic representation of the temporal sampling. In (a) Constant Count Input Mode (CCIM), the events are sampled from the continuous input event stream by splitting them into bins that include the same amount of events. In (b) Constant Time Input Mode (CTIM), each bin represents the same input duration. In both modes, every sample is then split into several subsets, based on the spatial coordinates of the events, as depicted in Figure 2. Note: blue and red points represent events with different polarities.
@@ 2.1.2 Normalized Local Patches to Tokens
The Feature Generator (FG) then converts each normalized input patch into an input token for the transformer model. The FG is shared between all the patches and consists in a customized shallow PointNet architecture. Every 4D active event contained in a single patch is processed through a shared MLP unit to generate a single high dimensional vector (of dimension $c$). The feature generator inherits the capacity of the PointNet to process an input of varying length. In our implementation, an MLP is composed of a succession of 1D-convolution layers.
Each 4D event is thus mapped to a high dimensional vector, and the Patch Feature is obtained by applying a channel-wise maximum operation on the resulting vectors. At this stage, all $P$ active patches have been converted into $P$ $c$-dimensional vectors, denoted as Patch Features (Fig. 5 (a)).
Normalizing and scaling the events with respect to the patch size results in information loss. In particular, the spatial origin of each local cloud in relation to the absolute (x,y) pixel grid is disregarded. Thus, the LERT module adds a positional embedding (i.e. a learnt linear embedding) to each Patch Feature, incorporating its spatial location within the predefined 2D grid. Positional embeddings supply the model with spatial distribution information about the tokens, they are learnt during training, and they are unique for every patch coordinate. A detailed schematic of the LERT embedding can be seen in Figure 4.
@@ 2.2.1 Introduction to the ALERT module
We thus introduce the Asynchronous LERT (ALERT) Module, which is used during inference. This variation is designed to efficiently update the token sequence as new events occur (asynchronous processing), making it particularly suited for online inference of event-based data. To achieve this, the ALERT module requires three key modifications during training and/or inference: handling continuous values of time, updating tokens with each new event, and memorizing past information.
@@ 2.2.2 Handling Continuous Time Values
The LERT module alters the time coordinates of events in a sequence, shifting each input stream so that the first event starts at time 00. This approach introduces inefficiencies in real-time asynchronous processing, as it hinders the feasibility of recurrent or iterative processing of events. Our proposed solution consists in encoding the time value by using a periodic function. We achieve this by representing time with two sinusoidal waves: $t_{x}(t)=\alpha\cos(2\pi ft+\phi)$, and $t_{y}(t)=\alpha\sin(2\pi ft+\phi)$; taking inspiration from the positional encoding originally used in Transformers. The amplitude ($\alpha$), frequency ($f$) and phase ($\phi$) of these waves are constant and tuned for each application. The Time Encoded (TE) LERT (TELERT) Module applies this periodical wrap prior to the PointNet model. So, TELERT must be the module used during training if ALERT is used for inference. The only change made before the channel-wise maximum operation between the LERT and TELERT/ALERT modules is this time encoding (see Figure 5 b). Next, we present a method applied at token stage that emulates a sliding buffer in input during prediction.
@@ 2.2.3 Asynchronous Event-Based Token Update
Add Information: The $k$ new event(s) are processed by TELERT. The resulting patch token is the channel-wise maximum between the previous patch token and the $k$ new event feature vector(s).
Forgetting information using Old Maximum Value Decay (OMVD): The “age” of the last update from each tokens’ channel is simply tracked by a counter. All channels whose associated counter channel is higher than a certain threshold undergo decay. All tokens are subject to this decay, not only the newly updated one, as the oldest information might be stored in different embeddings. The decay is applied in the form of an exponential (×$\times\text{e}^{-\lambda}$), where the decay rate $\lambda$ can be tuned for each model or scene characteristics, and could potentially be learned; the pseudocode can be found in Appendix A.1. This mechanism draws inspiration from the “leakage” concept in SNNs literature (Bouvier et al., 2019), but instead of applying it to all neurons/features of the model, we apply it only on tokens, to minimize compute.
@@ 2.2.4 Token Memorization and Activity Status
In ALERT, each new event triggers a corresponding patch token update. No raw event needs to be buffered. Instead, all tokens and associated age counter values are memorized throughout the entire runtime of the model, their patch being active or not. The non-activeness of the tokens is tracked using an event count for each patch. This counter is updated ($+1$ or $-1$) every time the token is updated (added information or OMVD, respectively).
@@ 2.3.1 Transformer Encoder for Classification
To evaluate the proposed LERT and ALERT modules, we plug them to a Vision Transformer model trained from scratch for classification tasks. As our focus is to assess the functionality of the proposed learnt synchronous and asynchronous embedding modules, we deploy a stack of standard transformer encoders, as in the original ViT (Dosovitskiy et al., 2021). The final classification head simply consists of a single linear layer followed by a SoftMax activation function.
@@ 2.3.2 Synchronous Training and Asynchronous Inference
During training and for synchronous inference (using LERT or TELERT), all samples are processed independently from each other. So, classic GPU/TPU accelerated processing is done as usual. Once TELERT is trained, the embedding module can be converted to ALERT for inference by adding the event update functionalities (adding information and OMVD).
The ALERT version, used only for inference, can process events continuously. However, applying the feature generator on every single event without distributed processing is extremely time consuming. Hence, for simulation purposes, the asynchronous inference pipeline is assessed by updating and processing the tokens on-demand every $\Delta t$. This has no impact on the resulting accuracy, as the mathematics of the asynchronous update are ultimately not altered. Proof can be found in Appendix A.2.